OpenStudy (anonymous):

Evaluate the following expression. 3^-2

Nnesha (nnesha):

\[\huge\rm x^{-m} = \frac{ 1 }{ x^m }\] exponent rule

OpenStudy (anonymous):

oh ya THX! :)

Nnesha (nnesha):

np :-)

OpenStudy (anonymous):

Evaluate the following expression. 70 What would i do here?

OpenStudy (anonymous):

7^0

Nnesha (nnesha):

anything to the 0 is just 1 \[\huge\rm (Nnesha)^0 = 1\] \[\huge\rm (PinHead_Larry99)^0 =1\] \[\huge\rm (4mn)^0 =1\] \[\huge\rm (chocolates)^0 =1\]

OpenStudy (anonymous):

oh wow LOL Thx again :)

Nnesha (nnesha):

ahah my pleasure :-)